# fMRI-3D: A Comprehensive Dataset for Enhancing fMRI-based 3D Reconstruction

### v1

@@ 6.1 Metrics
To effectively evaluate the performance of our models on this novel task, we employ metrics across two key dimensions: semantic-level and structure-level.
Semantic Level. To assess the semantic quality of our model, we adopt standard metrics commonly used in prior 2D fMRI studies [2, 3, 51, 52, 23], specifically N-way top-K accuracy. We report 2-way-top-1 and 10-way-top-1 accuracy, as shown in Table II. Additionally, we calculate the Learned Perceptual Image Patch Similarity (LPIPS) [53] to evaluate the perceptual quality of the semantic information. These metrics are computed by comparing the reconstructed images with the ground truth (GT) images. Both the reconstructed and GT objects are rendered into images at every 60-degree rotation, and the metrics are calculated for each frame. The final score is derived by averaging the values across frames.
Structure Level. In addition to semantic evaluation, it is essential to measure how well our model captures the geometrical structure of objects. We utilize common 3D reconstruction metrics [54, 55, 17]: Fréchet Point Cloud Distance (FPD, scaled by ), Chamfer Distance (CD, scaled by ), and Earth Mover’s Distance (EMD, scaled by ). These metrics are computed by sampling point clouds from the GT and reconstructed meshes. Additionally, we calculate the frame-averaged Structural Similarity Index (SSIM) using the same rendering process to evaluate similarity from various viewpoints.
@@ 6.2 Implementation Details
As detailed in Sec. 3, each data pair in the fMRI-Shape dataset consists of 10 fMRI frames and 192 corresponding images. For the experiment for fMRI-Shape, to maximize dataset usage and incorporate data augmentation, we randomly select 6 fMRI frames for training. The original 1023 × 2514 2D fMRI images are resized to 256 × 256 for processing. For the image data, 8 frames are randomly selected for training, with each image resized to 224 × 224. During inference, only the middle 6 fMRI frames are used, without additional images. Regarding the model architecture, we set to incorporate adapter layers.
The model training process is divided into two stages. In the first stage, we optimize the neuro-fusion encoder and the feature bridge diffusion model. The neuro-fusion encoder is initialized with pre-trained weights, while the feature bridge diffusion model is trained from scratch, which takes approximately 2.5 days on a single A100 GPU. In the second stage, all parameters, including those of the adapters, are optimized, while the parameters in Argus are frozen. This stage is completed in roughly one day on a single A100 GPU.
6.3 Standard Experiment
@@
As the first attempt to model 3D imaging within the human brain, we define the standard experiment setting as training and testing on the pre-split, person-specific Core set in fMRI-Shape, using the defined metrics to evaluate our model’s performance. Given the complexity of this task, which involves various brain regions, direct comparisons with existing models are not feasible. To reduce training costs, we adapted the LEA and fMRI-PTE models, which were trained on the same vision ROIs and have strong performance. These models utilize the same 3D decoder as MinD-3D. These adaptations serve as baselines in this study, enabling a more contextual and fair comparison within this novel domain.
Tab. II reports the averaged metrics at both the structural and semantic levels for all subjects. MinD-3D demonstrates superior performance across all metrics, achieving approximately 83.9% accuracy in 2-way-top-1, 43.2% accuracy in 10-way-top-1, and scores of 0.734, 0.230, 3.157, 1.742, and 3.833 in other metrics. These results indicate that our model not only excels in generating objects with high semantic accuracy but also performs exceptionally well in preserving structural similarity, outperforming baseline methods. Tab. II also highlights the effectiveness of both the diffusion model and the contrastive learning module, which play critical roles in the success of MinD-3D.
@@
Qualitative results from LEA-3D, fMRI-PTE-3D, and our method are presented in Fig. 8. MinD-3D consistently produces 3D objects that are structurally similar to their real counterparts, while maintaining semantic integrity in most cases. This highlights the robustness of our model in handling this challenging task and its ability to generate faithful reconstructions.
Furthermore, given the individual differences highlighted in Fig. 3, we also examine our model’s performance on each subject to assess its generalization ability. Fig. 9 shows several examples from 5 different subjects, demonstrating the effectiveness of the MinD-3D model. Due to varying brain responses among subjects to the same objects, some differences are observed in the results. However, overall, the model successfully reconstructs both the category and characteristics of the objects. This validates the feasibility of the proposed task, Recon3DMind, and the associated fMRI-Shape dataset developed for this study.
6.4 Out-Of-Distribution Experiments
@@
To effectively utilize a subset of the fMRI-Shape dataset and further assess the generalization capabilities of our proposed MinD-3D model, we conduct two Out-Of-Distribution (OOD) experiments under challenging settings: 1) Across-Person Testing (APT): In APT, we evaluate our model, which was trained only on Subject 1, using the data from Subject 9. We compare the results with the baselines and report the metrics in Tab. III. 2) Across-Person & Across-Class Testing (APACT): In APACT, we similarly evaluate our model, trained solely on Subject 1, with the data from Subject 11. We also compare with the baselines and report the metrics in Tab. III.
@@
We present the reconstructed objects from AP & APAC testing in Fig. 10. As shown in Fig. 3, individual differences significantly impact the results, which our AP & APAC tests empirically confirm. Despite the high difficulty, MinD-3D successfully recovers the basic shapes of the objects, providing a strong baseline for the community. While performance in these OOD scenarios does not reach In-Distribution (ID) levels—an expected outcome given the task’s complexity and the substantial individual differences and domain gaps—our method still surpasses existing baselines. This demonstrates the robustness of MinD-3D and establishes a new benchmark for future work.
| Methods | APT | APACT | ||||
@@
7 Analysis
7.1 Voxel Importance Analysis for Object Angle Variations
In this section, we explore the pattern when the angle of the object changes. Most directly, we perform linear regression on the whole-brain fMRI signal data for classification. The weight of each voxel in the brain serves as its importance score. Specifically, we performed logistic regression on the whole-brain fMRI signals using all voxels from subject 1 while viewing objects from two different angles (the 3rd and 9th frames of fMRI). We randomly selected 70% of the data for training and used the remaining 30% for testing. The accuracy of the linear classifier is 91.34%. To evaluate the importance of each voxel, we visualized the absolute values of the classifier weights in Fig. 11. Red regions indicate voxels with higher coefficients, while others represent lower coefficients. These ROIs include “2, AIP, FFC, FST, IPS1, LIPv, LO3, MIP, MST, MT, PH, PFt, PGp, PHT, TPOJ2, V1, V2, V3, V3A, V3B, V3CD, V4, V4t, V7, V8, VIP, 7PC, and 7PL.” These ROIs involve the parietal and occipital lobe, corresponding to spatial information processing and visual processing in the human brain, respectively.
@@ 7.2.1 Between Different Types of Objects
Building on the previous angle-based experiment, we perform logistic regression on the whole-brain fMRI signals of subject 1 to differentiate between cars and rifles. The linear classifier achieved an accuracy of 79.54%. We also visualize the absolute values of the classifier weights in Fig. 12, where red regions indicate voxels with higher coefficients and other colors represent lower coefficients. The regions of interest (ROIs) include “25, 47s, A5, AIP, PreS, STGa, STSda, TGd, V1, V2, V3, V3A, V4, V4t, V6, V7, and V8,” primarily located in the temporal and occipital lobes, corresponding to high-level visual functions and visual processing, respectively.
@@ 7.2.2 Between Objects Within the Same Category
In this case, classification experiments are more challenging due to the lack of specific labels for supervision. However, we visualize the absolute difference between two objects (a car and a plane) in Fig. 13 for analysis. In the figure, deep blue indicates voxels with higher values, while lighter colors represent lower values. Red rectangles highlight regions with higher values. Notably, these regions are predominantly located in the parietal and occipital lobes, which are known to play key roles in object differentiation.
7.3 Explore how our brain understand semantic information
Our proposed MinD-3D successfully reconstructs 3D objects both semantically and structurally, so we aim to explore how the brain processes semantic information of different objects. Our model mainly focuses on the visual regions and does not include a specific branch for semantic information. To achieve this, we use Class Activation Mapping (CAM) methods [56] to analyze the importance of each part of the input fMRI frame. We We present CAMs for three objects in Fig. 14. From this, we identify the following ROIs that may be involved in processing visual semantic information: FFC, FST, IP0, MT, MST, PHA1, PHA2, PHA3, PH, PIT, V1, V2, V3, V3A, V4, V4t, V6A, V7, V8, VMV1, VMV2, VMV3, and VVC.
7.4 Features Analysis
To understand the biological relevance of our feature extractor, in line with previous studies [19, 57, 58], we employ a linear encoding model to project the extracted fMRI features onto brain activity patterns, using subject 1 as a representative case for analysis. Specifically, we utilize ridge regression to establish the mapping between and the fMRI signals within the training dataset. We then project from the test set into the fMRI space and compute the Pearson correlation coefficient to assess the relationship. As illustrated in Fig. 15, areas positively correlated with visual Regions of Interest (ROIs) are marked in red, while other regions showing positive correlations are highlighted in blue. The results demonstrate a strong positive correlation between and visual ROIs, confirming the effective extraction of relevant features. Additionally, we perform a similar correlation analysis with the visual latent feature , which also shows positive correlations. Notably, the correlation strength of surpasses that of , indicating its superior relevance to brain activity. However, it is important to note that in some measures, the correlation coefficient of is higher than that of , which highlights the complexity of the task and underscores the critical role of bridging features. Furthermore, our results indicate that our generated visual latent features achieve activation levels comparable to the ground truth (GT), emphasizing the effectiveness of our comparative learning approach and demonstrating well-aligned fMRI feature and image spaces.

### v3

@@ 6 Experiments and Benchmark
Methods | Dataset | Semantic-Level | Structure-Level | Textural-Level | |||||
2-way | 10-way | FPD | CD | EMD | LPIPS | PSNR | SSIM | ||
LEA-3D [5] | fMRI-Shape | 0.787 | 0.371 | 4.229 | 2.291 | 5.347 | 0.557 | - | 0.617 |
fMRI-PTE-3D [4] | 0.815 | 0.392 | 3.571 | 1.992 | 4.621 | 0.462 | - | 0.645 | |
MinD-3D [15] | 0.828 | 0.459 | 3.157 | 1.742 | 3.833 | 0.306 | 14.98 | 0.674 | |
MinD-3D++ | 0.887 | 0.616 | 3.025 | 1.635 | 3.672 | 0.234 | 16.44 | 0.763 | |
Subject 1 | 0.912 | 0.645 | 2.953 | 1.531 | 3.419 | 0.226 | 17.22 | 0.783 | |
Subject 2 | 0.842 | 0.583 | 3.078 | 1.792 | 3.875 | 0.242 | 15.79 | 0.729 | |
Subject 3 | 0.869 | 0.623 | 3.043 | 1.642 | 3.773 | 0.239 | 15.98 | 0.751 | |
Subject 4 | 0.862 | 0.598 | 3.062 | 1.659 | 3.795 | 0.241 | 15.88 | 0.737 | |
Subject 5 | 0.908 | 0.638 | 2.985 | 1.567 | 3.511 | 0.228 | 17.16 | 0.785 | |
Subject 6 | 0.907 | 0.603 | 3.029 | 1.626 | 3.581 | 0.232 | 16.65 | 0.771 | |
Subject 7 | 0.894 | 0.613 | 3.034 | 1.635 | 3.672 | 0.235 | 16.02 | 0.768 | |
Subject 8 | fMRI-Shape | 0.901 | 0.627 | 3.018 | 1.525 | 3.498 | 0.232 | 16.79 | 0.778 |
MinD-3D [15] | 0.793 | 0.427 | 4.304 | 2.142 | 5.323 | 0.544 | 9.74 | 0.724 | |
MinD-3D++ | fMRI-Objaverse | 0.894 | 0.618 | 3.325 | 1.779 | 4.073 | 0.343 | 11.57 | 0.808 |
Subject 1 | 0.909 | 0.636 | 3.294 | 1.698 | 3.995 | 0.315 | 11.96 | 0.825 | |
Subject 6 | 0.893 | 0.616 | 3.311 | 1.751 | 4.032 | 0.326 | 11.67 | 0.811 | |
Subject 7 | 0.895 | 0.629 | 3.319 | 1.779 | 4.085 | 0.347 | 11.44 | 0.807 | |
Subject 8 | 0.882 | 0.593 | 3.353 | 1.863 | 4.134 | 0.367 | 11.20 | 0.797 | |
Subject 15 | fMRI-Objaverse | 0.891 | 0.616 | 3.347 | 1.802 | 4.117 | 0.358 | 11.23 | 0.801 |
@@ 6.1 Metrics
To effectively evaluate the performance of our models in reconstructing 3D objects from fMRI signals, we employ metrics across three primary dimensions: semantic, structural, and texture levels.
Semantic Level. To assess the semantic quality of our model, we use standard metrics commonly adopted in previous 2D fMRI studies [1, 2, 47, 48, 24], specifically N-way top-K accuracy. We report both 2-way top-1 and 10-way top-1 accuracies, as shown in Tab. II. These metrics are determined by comparing rendered images of the reconstructed objects with the ground truth (GT) images, which include texture.
Structural Level. Beyond semantic evaluation, it is crucial to measure how accurately our model captures the geometric structure of objects. We utilize common 3D reconstruction metrics [49, 50, 28]: Fréchet Point Cloud Distance (FPD) (scaled by ), Chamfer Distance (CD) (scaled by ), and Earth Mover’s Distance (EMD) (scaled by ). These metrics are computed by sampling point clouds from both the GT and the generated meshes.
Texture Level. To evaluate the quality of the texture and appearance in the reconstructed 3D objects, we use five metrics: Learned Perceptual Image Patch Similarity (LPIPS) [51], Structural Similarity Index (SSIM), and Peak Signal-to-Noise Ratio (PSNR), all calculated at the RGB pixel level.
All 3D objects are rendered in the same format as the input data for the multi-view diffusion model described in Sec. 5.3. These metrics are calculated for each frame, and the final scores are obtained by averaging the values across all frames.
@@ 6.2 Implementation Details
As detailed in Sec. 3, each sample in our dataset comprises 8 or 10 fMRI frames. To maximize dataset utilization and apply data augmentation during experiments, we randomly select 6 fMRI frames from each sample for training and use the middle 6 frames for inference. The vision region of interest (ROI), extracted from the original 2D fMRI images, is resized to for processing. During contrastive learning, we use the pretrained ViT-H-14 CLIP vision and text encoders to extract features from images and texts. For the image data, as discussed in Sec. 5.3, we use Blender to render all 3D objects into six distinct views and randomly select one image, resized to , for training. For textual descriptions, we use the category name of each object as the text input for the fMRI-Shape dataset. For the fMRI-Objaverse dataset, we adopt text descriptions sourced from Cap3D[52]. It is worth mentioning that our contrastive learning is not computed on the class token from the encoder’s final transformer layer but rather on results from an intermediate layer. Regarding the architecture, we configure LoRA with and within the multi-view diffusion model on the Q and V layers in the attention blocks, while keeping other parameters fixed during training. MinD-3D is trained end-to-end in a single stage, with both the encoder and the diffusion model initialized using pre-trained weights. Training each model takes approximately one day on eight A100 GPUs. We evaluate our model on both the fMRI-Shape and fMRI-Objaverse datasets.
6.3 Experiments on fMRI-Shape
6.3.1 Standard Experiment
@@
As the first effort to model 3D textured imaging within the human brain, we establish a standard experimental setup by training and testing on the pre-split, person-specific Core set in fMRI-Shape. Predefined metrics are used to evaluate the model’s performance. Given the task’s complexity—which involves multiple brain regions—direct comparisons with existing models are challenging, except with MinD-3D. To reduce training costs, we adapted the LEA and fMRI-PTE models, both trained on the same vision ROIs and demonstrating strong performance. Sharing the same 3D decoder as MinD-3D, these models serve as baselines in our experiments, enabling a more contextual and fair comparison within this novel domain.
@@
In the left part of Fig. 10, we present qualitative results of MinD-3D++ on the fMRI-Shape dataset. The model consistently generates 3D objects that are structurally and texturally similar to their real counterparts while maintaining semantic integrity in most cases. This underscores the robustness of our approach in handling a challenging task and its ability to produce faithful reconstructions. Importantly, the appearance of the reconstructed objects closely resembles the ground truth, demonstrating the effectiveness of our model.
Tab. II presents averaged metrics at three levels across all subjects alongside the baselines. MinD-3D++ outperforms MinD-3D and the other baselines on both semantic and structural levels, indicating its excellence in generating textured objects with high semantic accuracy and a strong ability to preserve structural similarity. Moreover, MinD-3D++, specifically designed to address the shortcomings of MinD-3D in terms of appearance and texture, significantly outperforms MinD-3D at the textural level.
Together, the qualitative and quantitative results validate the feasibility of reconstructing 3D textured objects from fMRI signals.
6.3.2 Out-Of-Distribution Experiments
@@
To effectively utilize a subset of the fMRI-Shape dataset and further assess the generalization capabilities of our proposed MinD-3D++ model, we conduct two Out-Of-Distribution (OOD) experiments under challenging settings: 1) Across-Person (AP) Testing: In AP testing, we evaluate our model, which was trained only on Subject 1, using the data from Subject 9. We compare the results with the baselines and report the metrics in Tab. III. 2) Across-Person & Across-Class (APAC) Testing: In APAC testing, we similarly evaluate our model, trained solely on Subject 1, with the data from Subject 11. We also compare with the baselines and report the metrics in Tab. III.
@@
We present the reconstructed objects from the AP and APAC tests in Fig. 11. As shown in Fig. 3, individual differences significantly impact the results, as confirmed by our empirical AP and APAC tests. Despite the high difficulty, MinD-3D++ successfully recovers the basic shapes of the objects, establishing a strong baseline for the community. We include the complete table of all metrics and subjects for the AP and APAC tests in the supplementary material. In Tab. 8, we compare the AP results with those from the regular experiment. While performance in these out-of-distribution (OOD) scenarios does not reach in-distribution (ID) levels—an expected outcome due to the task’s complexity and substantial individual differences and domain gaps—our method still outperforms existing baselines. This demonstrates the robustness of MinD-3D++ and establishes a new benchmark for future work.
Methods | AP | APAC | ||||
@@
MinD-3D++ | 3.582 | 2.237 | 4.975 | 5.347 | 3.912 | 6.915 |
6.4 Experiments on fMRI-Objaverse
The fMRI-Objaverse dataset presents a more challenging and larger-scale environment, increasing the difficulty of effective modeling. For this experiment, we randomly partitioned the objects into 2,709 for training and 432 for testing. Our goal was to reconstruct textured 3D objects from human brain data and rigorously evaluate our models’ capabilities. To this end, we trained MinD-3D++ on the pre-split fMRI-Objaverse dataset, using MinD-3D as our baseline for comparison.
As detailed in the lower section of Tab. II, we report metrics at three levels for our proposed model and all the baselines on the fMRI-Objaverse dataset. Across all metrics, MinD-3D++ outperforms MinD-3D and other baselines. Notably, on the more complex fMRI-Objaverse dataset—compared to the simpler fMRI-Shape—the performance improvement of MinD-3D++ is markedly greater, underscoring its robustness and enhanced capacity.
The right side of Fig. 10 presents qualitative results of MinD-3D++ on the fMRI-Objaverse dataset. Beyond capturing structural aspects comparable to MinD-3D, MinD-3D++ excels at accurately reconstructing appearance and color details. Furthermore, Fig. 12 offers additional 3D details. This figure displays the fMRI signals alongside corresponding RGB images, normals, depths, and masks, compared with ground truth (GT), to showcase the quality of the reconstructions results of MinD-3D++.
Given the substantially larger scale of the fMRI-Objaverse dataset relative to fMRI-Shape, the full capacity of MinD-3D++ is effectively leveraged. The high-quality textured 3D mesh reconstructions obtained from fMRI signals provide compelling evidence for the feasibility of reconstructing detailed textured 3D objects based on human brain data.
6.5 Ablation Study
To demonstrate the effectiveness of our proposed MinD-3D++, we conduct an ablation study on the contrastive loss. Specifically, we perform experiments in which we remove the contrastive loss, and separately remove the image and text branches from the contrastive loss, computing the loss based on the final transformer’s output from the encoder. We then report 2-way, 10-way, LPIPS, PSNR, and SSIM metrics compared with the full model in Tab. IV, which validates the effectiveness of incorporating image and text information and introducing contrastive learning at intermediate transformer layers.
Methods | 2-way | 10-way | LPIPS | PSNR | SSIM |
w/o contrastive | 0.830 | 0.463 | 0.296 | 15.17 | 0.697 |
w/o image | 0.843 | 0.549 | 0.266 | 15.79 | 0.712 |
w/o text | 0.858 | 0.563 | 0.252 | 15.93 | 0.734 |
Add on 12th | 0.875 | 0.604 | 0.248 | 16.24 | 0.759 |
Full model | 0.887 | 0.616 | 0.234 | 16.44 | 0.763 |
7 Analysis on Dataset
7.1 Analysis for Object Angle Variations
In this section, we explore brain activity patterns associated with object angle changes. Specifically, given whole-brain fMRI data collected while viewing 3D objects from different angles, we perform an angle classification task using simple linear classification on each subject in the fMRI-Shape dataset individually. We construct the fMRI data using views of objects from two different angles (the 3rd and 9th frames of each trial), drawn from both training and testing sets. We then randomly select 70% of the data for training and use the remaining 30% for testing. The average classification accuracy across all subjects for this angle-specific task reaches 88.10%.
To evaluate voxel importance, we visualize the absolute values of the classifier weights for Subject 1 in Fig. 13. Generally, the weight assigned to each voxel reflects its importance, with red regions indicating voxels with higher coefficients and other regions showing lower values. To better understand the relationship between ROIs and angle discrimination, we compute the mean value of the classifier weights within each HCP-MMP-defined ROI to quantify its importance. We then list the top-8 ROIs for each subject, along with the overall top-8 ROIs, calculated by summing the weights of each ROI across all subjects in the fMRI-Shape dataset, as shown in Tab. V. We also report the classification accuracy for each subject and the overall performance. Notably, ROIs such as V8, V3A, and V4 consistently appear among the top regions across all eight subjects, suggesting their common role in distinguishing object angles.
Furthermore, we conduct paired t-tests to evaluate whether these three ROIs are significantly more important than the primary visual area V1. The resulting p-values are , , and , respectively. All values fall below the significance threshold of , indicating that these ROIs are significantly more important than V1 in the context of angle discrimination.
@@ 7.2 Analysis of Different Objects
Subject | Accuracy | Top 8 Important ROIs |
1 | 91.34 | V3A, MT, V7, V8, V3, V3CD, V4, V3B |
2 | 84.88 | V8, V4, V3, V3A, VMV3, PHA1, MST, V1 |
3 | 88.82 | VMV3, V8, PIT, V3A, PEF, VVC, V4, FFC |
4 | 92.42 | MST, V3A, LO1, V8, V7, IPS1, V4, V4t |
5 | 83.07 | V8, V4, LO1, V3CD, V3A, V3, VMV3, V2 |
6 | 85.26 | V8, MST, VVC, MT, V4, FFC, 25, V3B |
7 | 92.40 | V3A, MST, LO1, V8, V7, V4t, V4, IPS1 |
8 | 86.59 | FEF, V6, VMV1, V8, VMV2, VVC, VMV3, 6v |
Average | 88.10 | V8, V3A, V4, MST, VMV3, LO1, V3, V7 |
Subject | Accuracy | Top 8 Important ROIs |
1 | 48.75 | LO1, PGp, V7, V3CD, V3, FFC, V4, V4t |
2 | 54.67 | PGp, V3CD, s32, 25, STGa, OFC, TA2, LO1 |
3 | 43.16 | LO1, FST, V4, V4t, s32, FFC, V6A, V3B |
4 | 48.22 | V8, PIT, V3, s32, V7, 25, FFC, V2 |
5 | 51.96 | LO1, 25, s32, V7, a24, V3CD, LO3, PGp |
6 | 43.54 | s32, MT, 10r, OFC, LO1, LO3, 25, V4 |
7 | 48.43 | MT, V4t, FST, V3, LO1, PH, V4, MST |
8 | 45.76 | LO1, 25, s32, V4, 52, STGa, V3CD, a24 |
Average | 48.10 | LO1, s32, 25, V4, PGp, V3, V7, V3CD |
@@ 7.2.1 Between Different Types of Objects
To observe and analyze how the brain responds to different object categories, we also perform a 13-class classification task by applying logistic regression to the whole-brain fMRI signals per subject in fMRI-Shape. The averaged accuracy of all the subjects achieved is 48.10%. Detailed classification accuracies for all subjects in the fMRI-Shape dataset, along with the top eight most important ROIs, are summarized in Tab. VI. The calculation of the top ROIs follows the same approach used for object angle variations (see Sec. 7.1).
We observe that ROIs such as LO1, s32, and area 25 consistently appear among the top regions across all eight subjects, suggesting their shared involvement in object recognition. Similarly, we conduct paired t-tests to evaluate whether these three ROIs are significantly more important than the primary visual area V1. The resulting p-values are , , and , respectively. The clear gap between these p-values and the significance threshold of further demonstrates the greater importance of these ROIs compared to V1 in the context of object recognition. Visualization of the absolute values of the classifier weights is given in Fig. 14, taking Subject 1 as an example. In this figure, red regions indicate voxels with higher coefficients; other colors show lower values.
Category | Top 8 Important ROIs |
Bench | PEF, LO2, 3b, 6v, FOP2, 6d, PHA2, 47m |
Cabinet | LO2, MST, IP1, 7AL, V4, TPOJ2, V3, TA2 |
Car | 5L, PFm, 6v, VMV3, FOP2, PF, 7AL, 4 |
Chair | 7AL, 7Am, PFm, 5L, 7PC, VIP, PGs, PGi |
Display | PFm, V4t, PGs, V3CD, LO2, VIP, 8Av, p9-46v |
Lamp | 7Am, 7AL, PFm, 7PC, 6ma, VIP, 6d, 1 |
Phone | 7AL, PFm, 7Am, LO2, 6mp, IP2, 1, 5L |
Plane | V4t, 6v, 6ma, 6d, 6a, 10d, 9p, PFm |
Rifle | PHA2, OP2-3, 10d, MST, TA2, V3B, OP4, PI |
Sofa | PFm, PGs, 7AL, 7Am, IP2, VIP, 5L, PEF |
Speaker | LO2, TPOJ2, V3B, V3A, 10d, V3, MST, TA2 |
Table | STSda, OP4, 43, FOP2, TA2, Ig, V6A, STSva |
Vessel | PFm, 7AL, 7Am, IP2, 6d, 7PC, 6mp, 6ma |
@@ 7.2.2 Between Objects Within the Same Category
In this case, classification experiments are more challenging due to the lack of specific labels for supervision. Therefore, we first explore the ROIs involved in recognizing intra-class objects for the same subject. As shown in Fig. 15, we visualize the absolute voxel-wise difference between two objects (a car and a plane) for analysis. In the figure, deep blue indicates voxels with higher values, while lighter colors represent lower values. Red rectangles highlight regions with particularly high values. This suggests that the ROIs required for recognizing intra-class objects within the same subject are correlated to some extent.
To further investigate this, we compute the mean absolute difference across all pairwise combinations of fMRI signals within the same object category, as shown in the following equation, to obtain the weight for each voxel:
We then average the voxel-wise weights within each ROI defined by the HCP-MMP atlas to quantify the importance of each region and select the top 8 ROIs based on their scores. We first take Subject 1 as an example to analyze the key ROIs involved in recognizing different intra-class objects. As shown in Tab. VII, we present the top 8 ROIs for Subject 1 when classifying intra-class objects across 13 different object categories in fMRI-Shape. Among them, regions such as PFm, 7AL, and 7Am consistently contribute to the recognition of most objects. The remaining top 8 ROIs for other subjects are in Tab. 1 to Tab. 7 in the supplementary material. To further examine whether consistent ROIs exist across subjects when recognizing intra-class objects, we compute the average weights across 13 categories for each subject using the same method applied to Subject 1. The top 8 ROIs for each subject across all categories are reported in Tab. VIII. We do not observe clear overlaps in these ROIs across subjects, which is understandable given the individual cognitive variability shown in Fig. 3.
Subject | Top 8 Important ROIs |
1 | PFm, 7AL, 7Am, LO2, PGs, IP2, 1, 7PC |
2 | 5L, 7AL, 7PL, VIP, 5m, 6ma, 7Am, 1 |
3 | PGs, 9a, 10d, PGp, p10p, 7PL, V6A, VIP |
4 | 9a, 9p, 5L, 8BL, 10d, V6A, 9m, LO2 |
5 | 5L, 10d, 7AL, 6v, 9a, 8BL, V6A, 5m |
6 | PF, IP1, PFt, PFm, IP2, 1, 7PC, PFop |
7 | 9a, 9p, 5L, 8BL, 10d, V6A, PFm, LO2 |
8 | V3CD, LO1, V3A, MT, LO2, V7, V3, PGp |
7.3 Explore how our brain understands semantics
MinD-3D++ successfully reconstructs 3D objects both semantically and structurally, so we aim to explore how the brain processes their semantic information. To this end, we use Class Activation Mapping (CAM)[53] to analyze the importance of each part of the input fMRI frame. CAMs for three objects are shown in Fig.16. From these, we identify the following ROIs potentially involved in visual semantic processing: FFC, FST, IP0, MT, MST, PHA1, PHA2, PHA3, PH, PIT, V1, V2, V3, V3A, V4, V4t, V6A, V7, V8, VMV1, VMV2, VMV3, and VVC.